# Why Does Dropping Edges Usually Outperform Adding Edges in Graph Contrastive Learning?

### v2

@@ Abstract
Graph contrastive learning (GCL) has been widely used as an effective self-supervised learning method for graph representation learning. However, how to apply adequate and stable graph augmentation to generating proper views for contrastive learning remains an essential problem. Dropping edges is a primary augmentation in GCL while adding edges is not a common method due to its unstable performance. To our best knowledge, there is no theoretical analysis to study why dropping edges usually outperforms adding edges. To answer this question, we introduce a new metric, namely Error Passing Rate (EPR), to quantify how a graph fits the network. Inspired by the theoretical conclusions, we propose a novel GCL algorithm, Error-PAssing-based Graph Contrastive Learning (EPAGCL), which uses both edge adding and edge dropping as its augmentation. To be specific, we generate views by adding and dropping edges according to the weights derived from EPR. Extensive experiments on various real-world datasets are conducted to validate the correctness of our theoretical analysis and the effectiveness of our proposed algorithm.
@@ 2.2 Adaptive Data Augmentations for Graph
Some researchers have investigated methods for adaptive data augmentations (Li 2022; Zhang et al. 2024) for GCL. JOAO (You et al. 2021) adaptively and dynamically selects data augmentations for specific graph data. GCA (Zhu et al. 2021) sets the probability of dropping edges and masking features according to their importance. ADGCL (Suresh et al. 2021) optimizes its adversarial graph augmentation strategies for contrastive learning. GCS (Wei et al. 2023) screens sub-structure in graphs with the help of a gradient-based graph contrastive saliency. And NCLA (Shen et al. 2023) learns graph augmentations by multi-head graph attention mechanism.
@@ 2.3 Adding Edges as Augmentation
Both methods tend to minimize the negative impact of the augmentation, which is the same as our proposed framework. To be specific, AdaEdge and GAUG reduce the probability of error augmentation, while EPAGCL reduces the magnitude of the impact if an error occurs.
@@ 3.3 Why is Adding Edges Usually Worse?
Under Assumption 3.4, if graph $\mathcal{G}$ is obtained through adding edge $(v_{i},v_{j})$ to $\mathcal{G}$, it holds that
{$r_{\mathcal{G}^{\prime}}-r_{\mathcal{G}}&=\frac{k}{m^{\prime}}\cdot\alpha_{i,j}&\text{if}\;c(v_{i})=c(v_{j}),\\ r_{\mathcal{G}^{\prime}}-r_{\mathcal{G}}&=\frac{1-k}{m^{\prime}}\cdot\alpha_{i,j}&\text{if}\;c(v_{i})\neq c(v_{j}),\end{aligned}\right.$

### v4

@@ Abstract
Graph contrastive learning (GCL) has been widely used as an effective self-supervised learning method for graph representation learning. However, how to apply adequate and stable graph augmentation to generating proper views for contrastive learning remains an essential problem. Dropping edges is a primary augmentation in GCL while adding edges is not a common method due to its unstable performance. To our best knowledge, there is no theoretical analysis to study why dropping edges usually outperforms adding edges. To answer this question, we introduce a new metric, namely Error Passing Rate (EPR), to quantify how a graph fits the network. Inspired by the theoretical conclusions and the idea of positive-incentive noise, we propose a novel GCL algorithm, Error-PAssing-based Graph Contrastive Learning (EPAGCL), which uses both edge adding and edge dropping as its augmentations. To be specific, we generate views by adding and dropping edges based on the weights derived from EPR. Extensive experiments on various real-world datasets are conducted to validate the correctness of our theoretical analysis and the effectiveness of our proposed algorithm. Our code is available at: https://github.com/hyzhang98/EPAGCL.
@@ 2.2 Adaptive Data Augmentations for Graph
Some researchers have investigated methods for adaptive data augmentations (Li 2022; Zhang et al. 2024) for GCL. JOAO (You et al. 2021) adaptively and dynamically selects data augmentations for specific data. GCA (Zhu et al. 2021) sets the probability of dropping edges and masking features according to their importance. ADGCL (Suresh et al. 2021) optimizes adversarial graph augmentation strategies for contrastive learning. GCS (Wei et al. 2023) screens sub-structure in graphs with the help of a gradient-based graph contrastive saliency. NCLA (Shen et al. 2023) learns graph augmentations by multi-head graph attention mechanism.
@@ 2.3 Adding Edges as Augmentation
Both methods tend to minimize the negative impact of the augmentation, which is the same as our proposed framework. To be specific, both of them reduce the probability of error augmentation, while EPAGCL, inspired by pi-noise (a theoretical framework to learn beneficial noise) (Li 2022), reduces the magnitude of the impact if an error occurs.
@@ 3.3 Why is Adding Edges Usually Worse?
Under Assumption 3.4, if graph $\mathcal{G}^{\prime}$ is obtained through adding edge $(v_{i},v_{j})$ to $\mathcal{G}$, it holds that
{$r_{\mathcal{G}^{\prime}}-r_{\mathcal{G}}&=-\frac{k}{m^{\prime}}\cdot\alpha_{i,j}&\text{if}\;c(v_{i})=c(v_{j}),\\ r_{\mathcal{G}^{\prime}}-r_{\mathcal{G}}&=\frac{1-k}{m^{\prime}}\cdot\alpha_{i,j}&\text{if}\;c(v_{i})\neq c(v_{j}),\end{aligned}\right.$